OpenStudy (devonhoward15):

Is anyone kind enough to stick with me... And help me with 4 big problems? work and everything

OpenStudy (kky10997):

ok

OpenStudy (devonhoward15):

you really good with algebra?

OpenStudy (devonhoward15):

@Vocaloid @agent0smith can you guys help me with 4 big problems?

Vocaloid (vocaloid):

sure

OpenStudy (devonhoward15):

thanks... if you want i can put each question in a individual question

Vocaloid (vocaloid):

sure

Vocaloid (vocaloid):

that works

OpenStudy (devonhoward15):

k... sorry its 3 questions one is 2 parts

OpenStudy (devonhoward15):

Solve each equation, if possible. Write irrational numbers in simplest radical form. Identify the strategy you used to get your solution. X^2-9=0

Vocaloid (vocaloid):

x^2 - 9 = 0 start by adding 9 to both sides

Vocaloid (vocaloid):

x^2 - 9 + 9 = ?

OpenStudy (devonhoward15):

x^2?

OpenStudy (devonhoward15):

x^2-18

Vocaloid (vocaloid):

good so x^2 = 9 take the square root of both sides to get x = 3 or x = -3 as your solutions

OpenStudy (devonhoward15):

i got it right the first time?

Vocaloid (vocaloid):

yes

Vocaloid (vocaloid):

for the 'strategy' name i would just say isolating the variable

OpenStudy (devonhoward15):

ok so x =-+3

Vocaloid (vocaloid):

good, what's next?

OpenStudy (devonhoward15):

let me write it down really quick

OpenStudy (devonhoward15):

ok for the one we just did i have to pick inverse operations, quadratic formula, factoring

Vocaloid (vocaloid):

inverse operations, then

Vocaloid (vocaloid):

(adding 9 was how we "inversed" the -9)

OpenStudy (devonhoward15):

x^2+4x+2=0

Vocaloid (vocaloid):

doesn't seem like we can factor it use the quadratic formula

Vocaloid (vocaloid):

a = 1, b = 4, c = 2

Vocaloid (vocaloid):

here it is if you need a refresher http://tothereal.files.wordpress.com/2013/05/quadratic-formula.jpg

OpenStudy (devonhoward15):

cant be solved

Vocaloid (vocaloid):

what do you mean, you just solved it

Vocaloid (vocaloid):

the answer exists, it's just not a whole number

OpenStudy (devonhoward15):

its an approximate number though? does that count as solved?

Vocaloid (vocaloid):

yes

Vocaloid (vocaloid):

the 2 root 2 part is exact

Vocaloid (vocaloid):

the only approximation is the decimal conversion

OpenStudy (devonhoward15):

ohh ok... now i gotta figure out how to do the work

Vocaloid (vocaloid):

just write down the formula show where you plugged in a, b, and c then just show the steps you used to simplify

OpenStudy (devonhoward15):

which is a,b,c?

Vocaloid (vocaloid):

a = 1, b = 4, c = 2

Vocaloid (vocaloid):

then plug them into the formula

OpenStudy (devonhoward15):

so it would be 1^2+4+2

Vocaloid (vocaloid):

what

Vocaloid (vocaloid):

write down the formula i gave you

Vocaloid (vocaloid):

now replace "a" with 1, "b" with 4, and "c" with 2 then simplify the expression

OpenStudy (devonhoward15):

ok

Vocaloid (vocaloid):

let me know what you get

Vocaloid (vocaloid):



Vocaloid (vocaloid):